# DCG-SQL: Enhancing In-Context Learning for Text-to-SQL with Deep Contextual Schema Link Graph

### v1

@@ Abstract
Text-to-SQL, which translates a natural language question into an SQL query, has advanced with in-context learning of Large Language Models (LLMs). However, existing methods show little improvement in performance compared to randomly chosen demonstrations, and significant performance drops when smaller LLMs (e.g., Llama 3.1-8B) are used. This indicates that these methods heavily rely on the intrinsic capabilities of hyper-scaled LLMs, rather than effectively retrieving useful demonstrations. In this paper, we propose a novel approach for effectively retrieving demonstrations and generating SQL queries. We construct a Deep Contextual Schema Link Graph, which contains key information and semantic relationship between a question and its database schema items. This graph-based structure enables effective representation of Text-to-SQL samples and retrieval of useful demonstrations for in-context learning. Experimental results on the Spider benchmark demonstrate the effectiveness of our approach, showing consistent improvements in SQL generation performance and efficiency across both hyper-scaled LLMs and small LLMs. Our code will be released.
@@ 3.2 Graph-based Demonstration Retrieval
To effectively train the graph representation model, we utilize contrastive learning, which ensures that similar graphs are clustered closely in the vector space, while dissimilar graphs are placed farther apart. We should identify positive demonstration samples that are useful for generating the SQL query for a given anchor. To achieve this, for the given anchor sample $x=(G_{x},SQL_{x})$, we first extract likelihoods of the $SQL_{x}$ when other samples $S=\{s_{i}=(G_{i},SQL_{i})\mid i=1,2,\dots\}$ are provided as demonstrations using Llama 3.1-8B-Instruct model Dubey et al. (2024). The likelihood is as follows:
@@
To identify positive and negative samples, we rank the $s_{i}$’s based on their likelihood of generating the anchor’s target SQL, treating high-ranked ones as positives and low-ranked ones as negatives. Since the retrieval model is trained to increase similarity with high-ranked demonstrations and decrease similarity with low-ranked ones, we can retrieve useful demonstrations from the retrieval pool through similarity of the graph embeddings Karpukhin et al. (2020); Rubin et al. (2022); Li et al. (2023c).
During inference, we compare the embedding of user input graph $G_{user}$ generated by the encoder, with the embeddings of the external DB pool $S(G,SQL)$ by computing the similarity. Finally, we select the top-k samples with the highest similarity as demonstrations.
@@ 3.3 SQL generation with In-context Learning
With the retrieved useful demonstrations, LLM can generate target SQL query correctly. In addition, we utilized automated-CoT prompting method to guide the SQL generation more effectively Zhang et al. (2022, 2023a). The retrieved demonstrations are categorized into types such as simple, join, and nested queries. A predefined CoT template is used for each type, with its content dynamically adapted to each demonstration to reflect the SQL generation process. This approach ensures that the model follows more consistent and structured steps when generating SQL queries. Details are provided in Appendix C.2.
@@ 4.1 Implementation Details
All experiments are conducted using the PyTorch framework on an A100 80GB GPU. Our method requires only small size of models for efficiency. For schema link graph construction, we train the schema pruning model based on RoBERTa-large, which has approximately 355 million parameters Liu et al. (2019). To link schema with tables or columns, we set $\tau_{tab}$ and $\tau_{col}$ to 0.66 and 0.43, respectively. For graph-based demonstration retrieval, We train a graph encoder model based on a bi-encoder architecture Karpukhin et al. (2020), where each encoder follows the Relation-Aware Transformer Shaw et al. (2018) structure initialized with BERT-base Devlin et al. (2019), totaling approximately 220 million parameters.
To ensure a fair comparison, we reproduce experimental results or follow existing studies. For SQL generation, we use various hyper-scaled LLMs and small LLMs, including GPT-4, GPT-3.5-Turbo, DeepSeek-Coder-33B-Instruct, Llama 3.1-8B-Instruct, DeepSeek-Coder-6.7B-Instruct, and Llama 3.2-3B-Instruct OpenAI et al. (2024); DeepSeek-AI et al. (2024); Dubey et al. (2024). In all experiments, the generation temperature is set to 0.5 demonstrations in our few-shot setting and our input prompt is introduced in Appendix C.1. Training the pruning model took approximately 3 hours, while the graph encoder model was trained for 8 hours. We used spaCy Honnibal and Montani (2017) and NLTK Bird and Loper (2004) libraries for the edges between question tokens.
@@ 4.5 Ablation Studies
Effectiveness of our retrieval method.
The Table 4 presents the results when our retrieval method is not applied and demonstrations are randomly chosen. Existing approaches Zhang et al. (2023a); Gao et al. (2024) failed to maximize the capabilities of in-context learning due to their ineffective retrieval approaches as shown in Table 1, whereas our method consistently achieves performance improvements across various datasets and LLMs.
To demonstrate the effectiveness of our retrieval method, we compare it with various retrieval approaches, as shown in Table 5. For demonstration retrieval, case (1) is based on the text embedding of question and (2) is based on the masked question text embedding Guo et al. (2024). Case (3) is based on both question and DB schema, but uses text embeddings. In case (4), it is based on the existing traditional graph representation Li et al. (2023b). In our case, it is based on deep contextual schema link graph. Cases (1)–(3) use BERT-base as the text encoder, whereas case (4) employs the same graph encoder as our method. All (1)–(4) are trained using the same self-supervised learning approach described in Section 3.2
As shown in cases (1) and (2), demonstration retrieval based only on question information does not achieve good performance. Also, retrieval methods that consider both question and DB schema information, but do not account for their contextual relationships, show poor performance. Our method effectively capture the relationship for the joint representation of Text-to-SQL samples. These results demonstrate that our deep contextual schema link-based graph retrieval effectively utilizes the DB schema information, which must be considered when generating SQL queries.
Effectiveness of our DB schema pruning.
Transferability of our proposed modules.
To demonstrate the transferability of our modules, we applied our retrieval to the DAIL-SQL Gao et al. (2024). All the experimental results in Table 7 are obtained using the same setting as DAIL-SQL. When we use the demonstrations sets retrieved by our method, the performance is 75.6% in EX and 61.3% in EM, respectively. This corresponds to improvements of 6.2% and 11.9% over the original DAIL-SQL. These results demonstrate that our contextual schema link-based retrieval method can be effectively applied to various prompt templates and Text-to-SQL methods.
Analysis of retrieved demonstrations.
To analyze the quality of retrieved demonstrations, we calculate the average Tree Edit Distance (TED) Yu et al. (2018a), which measures the similarity between the target SQL of user input and the SQLs of the retrieved demonstrations. If the TED is small, it indicates that the SQLs of retrieved demonstrations are similar to the target SQL of user input. As shown in the Table 8, our method achieves the lowest TED compared to existing approaches. Since in-context learning of LLMs tends to follow the patterns in the demonstrations Brown et al. (2020); Luo et al. (2024), the result supports that our proposed method effectively retrieves useful demonstrations for generating correct SQL.

### v2

@@ Abstract
Text-to-SQL, which translates a natural language question into an SQL query, has advanced with in-context learning of Large Language Models (LLMs). However, existing methods show little improvement in performance compared to randomly chosen demonstrations, and significant performance drops when smaller LLMs (e.g., Llama 3.1-8B) are used. This indicates that these methods heavily rely on the intrinsic capabilities of hyper-scaled LLMs, rather than effectively retrieving useful demonstrations. In this paper, we propose a novel approach for effectively retrieving demonstrations and generating SQL queries. We construct a Deep Contextual Schema Link Graph, which contains key information and semantic relationship between a question and its database schema items. This graph-based structure enables effective representation of Text-to-SQL samples and retrieval of useful demonstrations for in-context learning. Experimental results on the Spider benchmark demonstrate the effectiveness of our approach, showing consistent improvements in SQL generation performance and efficiency across both hyper-scaled LLMs and small LLMs. The code is available at https://github.com/jjklle/DCG-SQL.
@@ 3.2 Graph-based Demonstration Retrieval
To effectively train the graph encoder, we utilize contrastive learning that clusters similar graphs closely in the vector space while distancing dissimilar ones farther apart. We define positive demonstration samples as those that are useful for generating the SQL query of a given anchor. To achieve this, for the given anchor sample $x=(G_{x},SQL_{x})$, we extract likelihoods of the $SQL_{x}$ when other samples $S=\{s_{i}=(G_{i},SQL_{i})\mid i=1,2,\dots\}$ are provided as demonstrations using Llama 3.1-8B model Dubey et al. (2024). The likelihood is as follows:
@@
To identify positive and negative samples, we rank $s_{i}$’s based on their likelihood of generating the anchor’s target SQL, treating high-ranked ones as positives and low-ranked ones as negatives. Since the retrieval model is trained to increase similarity with high-ranked demonstrations and decrease similarity with low-ranked ones, we can retrieve useful demonstrations from the retrieval pool through similarity of the graph embeddings Karpukhin et al. (2020); Li et al. (2023c).
However, computing the likelihood of $SQL_{x}$ over the entire training dataset as demonstrations is computationally expensive. To address this, we construct a reduced candidate set for each anchor by selecting 100 structurally similar SQLs, using a tree edit distance computed over abstract syntax trees Yu et al. (2018a). Unlike Shen et al. (2024), which selects demonstrations based on AST similarity, we further re-rank the candidates by their usefulness for generating the anchor SQL. This enables us to identify samples that appear structurally similar but are unhelpful as demonstrations—performing hard negative mining.
During inference, we compare the embedding of user input graph $G_{user}$ generated by the encoder, with the embeddings of the demonstration pool by computing the similarity. We select top-k samples with the highest similarity as demonstrations.
@@ 3.3 SQL generation with In-context Learning
With the retrieved demonstrations, LLM can generate target SQL query correctly. In addition, we utilized automated-CoT prompting method to guide the SQL generation more effectively Zhang et al. (2022, 2023a). The retrieved demonstrations are categorized into types such as simple, join, and nested queries. Each type uses a predefined CoT template, dynamically adapted per demonstration to guide consistent and structured SQL generation. Details are provided in Appendix C.2.
@@ 4.1 Implementation Details
All experiments are conducted on an A100 80GB GPU. Our method requires only small size of models for efficiency. For schema link graph construction, we train the schema pruning model based on RoBERTa-large, which has approximately 355 million parameters Liu et al. (2019). For schema linking, we set $\tau_{tab}$ and $\tau_{col}$ to 0.66 and 0.43, respectively. For graph-based demonstration retrieval, We train a graph encoder model based on a bi-encoder architecture Karpukhin et al. (2020), where each encoder follows the Relation-Aware Transformer Shaw et al. (2018) structure initialized with BERT-base Devlin et al. (2019), totaling approximately 220 million parameters.
For a fair comparison, we reproduce experimental results or follow existing studies. For SQL generation, we use various LLMs, including GPT-4, GPT-3.5-Turbo, DeepSeek-Coder-33B-Instruct, Llama 3.1-8B-Instruct, DeepSeek-Coder-6.7B-Instruct, and Llama 3.2-3B-Instruct OpenAI et al. (2024); DeepSeek-AI et al. (2024); Dubey et al. (2024). In all experiments, the generation temperature is set to 0. We set 5 demonstrations for the few-shot setting and our input prompt is introduced in Appendix C.1. Training the pruning model took approximately 3 hours, while the graph encoder model was trained for 8 hours. We used spaCy Honnibal and Montani (2017) and NLTK Bird and Loper (2004) libraries for the edges between question tokens.
@@ 4.5 Ablation Studies
Effectiveness of our retrieval method.
Table 4 presents the results when demonstrations are randomly chosen. Existing approaches Zhang et al. (2023a); Gao et al. (2024) failed to maximize in-context learning capabilities due to their ineffective retrieval approaches as shown in Table 1, whereas our method consistently achieves performance improvements across various datasets and LLMs.
To demonstrate the effectiveness of our retrieval method, we compare it with various retrieval approaches, as shown in Table 5. For demonstration retrieval, case (1) is based on the text embedding of question and (2) is based on the masked question text embedding Guo et al. (2024). Case (3) is based on both question and DB schema, but uses text embeddings. In case (4), it is based on the existing graph representation Li et al. (2023b). Cases (1)–(3) use BERT-base as the text encoder, whereas case (4) employs the same graph encoder as our method. All (1)–(4) are trained using the same self-supervised learning approach described in Section 3.2
As shown in cases (1) and (2), demonstration retrieval based only on question does not achieve good performance. Also, retrieval methods that consider both question and DB schema information without their contextual relationships, show poor performance. In contrast, our method captures contextual relationships in the joint representation of Text-to-SQL samples, leading to better demonstration retrieval.
Effectiveness of our DB schema pruning.
Transferability of our proposed modules.
To demonstrate the transferability of our modules, we applied our retrieval to the DAIL-SQL Gao et al. (2024). All the experimental results in Table 7 are obtained using the same setting as DAIL-SQL. When we use the demonstrations sets retrieved by our method, the performance is 75.6% in EX and 61.3% in EM, respectively. This corresponds to improvements of 6.2% and 11.9% over the original DAIL-SQL. These results demonstrate that our retrieval method can be effectively applied to various prompt templates and Text-to-SQL methods.
Analysis of retrieved demonstrations.
To analyze the quality of retrieved demonstrations, we calculate the average Tree Edit Distance (TED) Yu et al. (2018a), which measures the similarity between the target SQL of user input and the SQLs of the retrieved demonstrations. If TED is small, it indicates that the SQLs of retrieved demonstrations are similar to the target SQL of user input. As shown in the Table 8, our method achieves the lowest TED compared to existing approaches. Since in-context learning of LLMs tends to follow the patterns in the demonstrations Brown et al. (2020); Luo et al. (2024), the result supports that our proposed method effectively retrieves useful demonstrations for generating correct SQL.